A baker made thrice as many cakes as pies. He made 50 more cakes than pies. How many cakes did he make
Answer by LinnW(1048)

let c = number of cakes
p = number of pies
c = 3p
c = p + 50
Set the two right hand sides to be equal
3p = p + 50
Subtract p from each side.
2p = 50
So
p = 25
Substitute in c = p + 50
c = 25 + 50
c = 75